# Inflation in light of ACT/SPT: a new perspective from Weyl gravity

### v1

@@ Abstract
Recent measurements from the Atacama Cosmology Telescope (ACT) and the South Pole Telescope (SPT) have placed the strictest constraints on the primordial scalar perturbation spectrum, reporting a spectral index of $n_{s}\sim 0.967-0.98$ at 95% confidence level. This result indicates a stronger scale invariance of the scalar perturbation than earlier estimates, posing challenges for numerous inflation models. In this work, we propose an appealing inflationary scenario from the Weyl scale-invariant gravity theory dominated by the higher-order curvatures. Specifically, the exponential curvature extensions are introduced to suppress the mass divergence of the inflaton. We find such scenario naturally yields leading-order predictions of $n_{s}\simeq 1-3/(2N)\sim 0.97-0.975$ or $n_{s}\simeq 1-5/(3N)\sim 0.967-0.972$ for various models, in excellent agreement with the ACT/SPT constraints. This result builds a concrete bridge between theoretical and observational scale invariance, implying an enduring cosmic echo of the primordial symmetry.
@@
Introduction — It is well established that the observations of cosmic microwave background (CMB) reveal an almost scale-invariant scalar perturbation spectrum, characterized by a spectral index $n_{s}$ slightly less than 1. Recently, the Atacama Cosmology Telescope (ACT) [1, 2] and the South Pole Telescope (SPT) [3] have released updated CMB measurements, indicating an $n_{s}$ value closer to 1 than previously reported in Planck 2018 [4]. Specifically, $n_{s}\sim 0.967-0.98$ in the 95% confidence interval. This suggests that the scale invariance of the scalar spectrum is even stronger than earlier estimates.
@@
An important goal of modern cosmology is to understand the origin of this approximate scale invariance. Generally, inflation stands as the leading theoretical framework capable of generating such a nearly scale-invariant spectrum [5]. Over the years, numerous inflation models have been developed. However, with the increasing precision of CMB observations—particularly following the release of the ACT/SPT results—many of these models are now facing stringent tests. For instance, well-studied scenarios such as Starobinsky inflation [6], Higgs inflation [7], and $\alpha$-attractors [8] predict a value of $n_{s}$ that is now too low to be consistent with the latest CMB data, leaving them disfavored at the 95% confidence level. These developments urge a reassessment of the inflationary mechanism, especially its intrinsic connection to scale invariance.
@@
An enlightening idea posits that inflation could emerge from a gravity theory with local scale symmetry, also known as Weyl scale invariance. This symmetry is first introduced by H. Weyl [9, 10], and now has been widely applied in particle physics[11, 12, 13, 14, 15, 16, 17], cosmology[18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], and quantum gravity [45, 46]. In such a framework of inflation, the spontaneous breaking of the Weyl symmetry would naturally give rise to a nearly scale-invariant scalar spectrum, implying a profound connection between the theoretical and observational scale invariance.
@@
In previous studies, a scale-invariant linear curvature term $\phi^{2}\hat{R}$ is often introduced alongside the quadratic term to realize the above scenario [18, 19, 20, 21, 22]. This model leads to a Higgs-like potential for inflation. However, its predicted value for $n_{s}$ is lower than that of the Starobinsky model, thus further exacerbating the tension with current observational constraint. In this work, we explore the role of higher-order curvature extensions—rather than a linear term—in departing the pure de Sitter state derived from $\hat{R}^{2}$ model and explaining the tiny deviation of scalar spectrum from scale invariance. Specifically, we consider exponential curvature extensions, which, by contrast with simple polynomial terms, significantly alleviate the mass divergence of inflaton. We calculate the cosmological observables for such models both analytically and numerically, confirming their consistency with the ACT/SPT results.
@@
Now we translate the Lagrangian to a more familiar form. Compared to the conventional Ricci scalar $R$, one can easily verify that
@@
Models — In this framework, the curvature term $F(\hat{R},\phi)$ determines the configuration of the effective potential $V(\Phi)$, which can be derived from Eqs. (7), (9), and (13), albeit the analytic expression may be unavailable for some complex cases. With these equations, one can readily verify the conclusions we mentioned earlier: the purely geometric case $F=\alpha\hat{R}^{2}$ is equivalent to a cosmological constant, and an additional scalar $\phi$ is able to make the model dynamical, leading to a slight deviation from the scale-invariant spectrum.
@@
Note that the simplification of exponential-extended models with Eq. (16) is only valid in the regime $\phi^{2}\gg R$. The behavior near $\phi=0$, particularly the divergence behavior, is presented in different manners. Since Eq. (7) in these models lacks explicit solutions, we construct an asymptotic solution in the limit $\phi\to 0$. To leading-order, all models yield the same estimate: $\chi_{\phi\to 0}\simeq 2\phi^{2}\ln\left|\frac{M_{P}}{\phi}\right|$. Substituting this into Eq. (9) gives
@@
The second derivative is still divergent, but the degree of divergence has reduced to the logarithmic level. This means the invalid interval of $\Phi$ where $m_{\Phi}>M_{P}$ becomes extremely narrow for sufficiently large $\zeta$. Specifically, it can be estimated to $|\Phi|<e^{-\zeta/2}$ for $\zeta\gg 1$. In fact, $\zeta$ allows for a very large value up to $10^{10}$ (shown in the next section), thus the premise of our discussion is satisfied.
@@
$\epsilon_{V}=2\xi\left(u^{2}-1\right)\left(\frac{u-1}{2u+1}\right)^{2},$ (28)
@@
$r\simeq\frac{\sqrt{3}}{3\sqrt{\xi}N^{3/2}}+\frac{23}{48\xi N^{2}}.$ (33)
@@
Discussion — A comparison with conventional $f(R)$ theories is instructive. In typical $f(R)$ inflation, the higher-order extensions are treated as corrections to the Starobinsky model [47, 48, 49, 50, 51, 52, 53, 54, 55]. From the perspective of phenomenology, the Starobinsky prediction, $n_{s}\simeq 1-2/N\sim 0.96-0.967$, lies slightly below the ACT/SPT preferred range. Positive higher-order curvature corrections would further diminish $n_{s}$, exacerbating the tension with observations. Conversely, negative corrections could raise $n_{s}$ into agreement. However, the parameter fine-tuning is required to some extent. Our models are in a different situation. The leading-order result of $n_{s}$ conforms directly to the ACT/SPT constraints, thereby avoiding fine-tuning. Moreover, from the perspective of theoretical completeness, inflation in the $f(R)$ framework with negative corrections no longer originates from a de Sitter state with constant curvature. The Ricci scalar $R$ is not convergent toward the beginning of inflation, and a critical value of $R_{c}$ (or corresponding $\Phi_{c}$ in the Einstein frame) emerges where $f(R)$ or $f^{\prime\prime}(R)$ turns negative. This necessitates a cut-off at $R_{c}$, rendering the theory UV incomplete [47, 54]. In contrast, our models preserve the UV completeness, which can be verified that $R_{\phi\to\infty}=M_{P}^{2}/(2\alpha)$, corresponding to a constant-curvature origin.
@@
Conclusion — In this work, we demonstrate a viable scenario that inflation arises from the higher-order curvature extensions, especially the divergence-suppressed extensions with exponential forms, on the basis of the Weyl scale-invariant $\hat{R}^{2}$ model. In these models, a ”pit” is developed on the cosmological constant potential in the Einstein frame, thereby deviating the scalar perturbation spectrum from exact scale invariance. Remarkably, such deviations align perfectly with the latest CMB observations across a broad parameter range. At the leading-order, the spectral index is given by $n_{s}\simeq 1-3/(2N)$ or $n_{s}\simeq 1-5/(3N)$ for various models, which lies comfortably within the 95% confidence interval reported by ACT/SPT. Consequently, a scale-invariant inflationary framework naturally explains the observed nearly scale-invariant spectrum, implying an enduring cosmic echo of the primordial symmetry.

### v2

@@ Abstract
Recent measurements from the Atacama Cosmology Telescope (ACT) and the South Pole Telescope (SPT) have placed the strictest constraints on the primordial scalar perturbation spectrum, reporting a spectral index of $n_{s}\sim 0.967-0.98$ at 95% confidence level. This result indicates a stronger scale invariance of the scalar perturbation than earlier estimates, posing challenges for numerous inflation models. In this work, we propose a novel inflationary scenario based on Weyl scale-invariant gravity, where the quadratic curvature establishes the scale invariance of the scalar spectrum, while the higher-order extensions imprint the observed slight deviation. Specifically, the exponential curvature extensions are introduced to suppress the mass divergence of the inflaton. We find such scenario naturally yields leading-order predictions of $n_{s}\simeq 1-3/(2N)\sim 0.97-0.975$ or $n_{s}\simeq 1-5/(3N)\sim 0.967-0.972$ for various models, in excellent agreement with the ACT/SPT constraints. This result builds a concrete bridge between theoretical and observational scale invariance, implying an enduring cosmic echo of the primordial symmetry.
@@
Introduction — It is well established that the observations of cosmic microwave background (CMB) reveal an almost scale-invariant scalar perturbation spectrum, characterized by a spectral index $n_{s}$ slightly less than 1. Recently, the Atacama Cosmology Telescope (ACT) [1, 2] and the South Pole Telescope (SPT) [3] have released updated CMB measurements, indicating an $n_{s}$ value closer to 1 than previously reported in Planck 2018 [4], especially after incorporating the latest baryon acoustic oscillation data from DESI [5]. Specifically, they constrain $n_{s}\sim 0.967-0.98$ in the 95% confidence interval. This suggests that the scale invariance of the scalar spectrum is even stronger than earlier estimates.
@@
An important goal of modern cosmology is to understand the origin of this approximate scale invariance. Generally, inflation stands as a compelling theoretical framework capable of generating such a nearly scale-invariant spectrum [6]. Over the years, numerous inflation models have been developed. However, with the increasing precision of CMB observations—particularly following the release of the ACT/SPT results—many of these models are now facing stringent tests. For instance, well-studied scenarios such as Starobinsky inflation [7], Higgs inflation [8], and $\alpha$-attractor T-models [9] predict a value of $n_{s}$ that is now too low to be consistent with the latest CMB data, leaving them disfavored at the 95% confidence level. These developments urge a reassessment of the inflationary mechanism, especially its intrinsic connection to scale invariance.
@@
An enlightening idea posits that inflation could emerge from a gravity theory with local scale symmetry, also known as Weyl scale invariance. This symmetry is first introduced by H. Weyl [10, 11], and now has been widely applied in particle physics [12, 13, 14, 15, 16, 17, 18], cosmology [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], and quantum gravity [47, 48]. In such a framework of inflation, the spontaneous breaking of the Weyl symmetry would naturally give rise to a nearly scale-invariant scalar spectrum, implying a profound connection between the theoretical and observational scale invariance.
@@
In previous studies, a scale-invariant linear curvature term $\phi^{2}\hat{R}$ is often introduced alongside the quadratic term to realize the above scenario [19, 20, 21, 22, 23]. This model leads to a Higgs-like potential for inflation. However, its predicted value for $n_{s}$ is lower than that of the Starobinsky model, thus further exacerbating the tension with current observational constraint.
In this work, we explore the role of higher-order curvature extensions—rather than a linear term—in departing the pure de Sitter state derived from $\hat{R}^{2}$ model. Such extensions are not simple corrections to prior models, but rather lead to a brand new inflationary scenario. Specifically, we consider exponential curvature extensions, which, by contrast with simple polynomial terms, significantly alleviate the mass divergence of inflaton. We calculate cosmological observables for such models and confirm their consistency with the ACT/SPT results.
@@
Now we translate the Lagrangian to a more familiar form. Compared to the conventional Ricci scalar $R$, one can verify that
@@
Models — In this framework, $F(\hat{R},\phi)$ determines the configuration of the effective potential $V(\Phi)$, which can be derived from Eqs. (5), (7), and (11), albeit the analytic expression may be unavailable for some complex cases. With these equations, one can readily verify the conclusions we mentioned earlier: the purely geometric case $F=\alpha\hat{R}^{2}$ is equivalent to a cosmological constant, and an additional scalar $\phi$ is able to make the model dynamical, leading to an inflation model.
@@
Note that the simplification of exponential-extended models with Eq. (14) is only valid in the regime $\phi^{2}\gg\hat{R}$. The behavior near $\phi=0$, particularly the divergence behavior, is presented in different manners. Since Eq. (5) in these models lacks explicit solutions, we construct an asymptotic solution in the limit $\phi\to 0$. To leading-order, all models yield the same estimate: $\chi_{\phi\to 0}\simeq 2\phi^{2}\ln\left|\frac{M_{P}}{\phi}\right|$. Substituting this into Eq. (7) gives
@@
The second derivative is still divergent, but the degree of divergence has reduced to the logarithmic level. This means the invalid interval of $\Phi$ where $m_{\Phi}>M_{P}$ becomes extremely narrow for sufficiently large $\zeta$. Specifically, it can be estimated to $|\Phi|<e^{-\zeta/2}M_{P}$ for $\zeta\gg 1$. In fact, $\zeta$ allows for a very large value up to $10^{10}$ (shown in the next section), thus the premise of our discussion is satisfied.
@@
$\epsilon_{V}=2\xi\frac{\left(u^{2}-1\right)\left(u-1\right)^{2}}{\left(2u+1\right)^{2}},$ (26)
@@
$r$ $\simeq\frac{1}{\sqrt{3\xi}N^{3/2}}+\frac{23}{48\xi N^{2}}.$ (31)
@@
Discussion — A comparison with conventional $f(R)$ theories is instructive. In typical $f(R)$ inflation, the higher-order extensions are treated as corrections to the Starobinsky model [49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. From the perspective of phenomenology, the Starobinsky prediction, $n_{s}\simeq 1-2/N\sim 0.96-0.967$, lies slightly below the ACT/SPT preferred range. Positive higher-order curvature corrections would further diminish $n_{s}$, exacerbating the tension with observations. Conversely, negative corrections could raise $n_{s}$ into agreement. However, the parameter fine-tuning is required to some extent. Our models are in a different situation. The leading-order result of $n_{s}$ conforms directly to the ACT/SPT constraints, thereby avoiding fine-tuning. Moreover, from the perspective of theoretical completeness, inflation in the $f(R)$ framework with negative corrections no longer originates from a de Sitter state with constant curvature. The Ricci scalar $R$ is not convergent toward the beginning of inflation, and a critical value of $R_{c}$ (or corresponding $\Phi_{c}$ in the Einstein frame) emerges where $f(R)$ or $f^{\prime\prime}(R)$ turns negative. This necessitates a cut-off at $R_{c}$, rendering the theory UV incomplete [49, 57]. In contrast, our models preserve the UV completeness, which can be verified that $R_{\phi\to\infty}=M_{P}^{2}/(2\alpha)$, corresponding to a constant-curvature origin.
@@
Conclusion — In this work, we demonstrate a viable scenario that inflation arises from the higher-order curvature extensions, especially the divergence-suppressed extensions with exponential forms, on the basis of the Weyl scale-invariant $\hat{R}^{2}$ model. In these models, the $\hat{R}^{2}$-induced cosmological constant serves as the origin of scale invariance for the scalar perturbation spectrum, while the higher-order curvatures bring about the observed slight deviation. Remarkably, such deviations align perfectly with the latest CMB observations across a broad parameter range. At the leading-order, the spectral index is given by $n_{s}\simeq 1-3/(2N)$ or $n_{s}\simeq 1-5/(3N)$ for various models, which lies comfortably within the 95% confidence interval reported by ACT/SPT. Consequently, a scale-invariant inflationary framework naturally explains the observed nearly scale-invariant spectrum, implying an enduring cosmic echo of the primordial symmetry.